An urn contains 6 white, 4 red and 9 black balls. If 3 balls are drawn at random, find
the probability that: (i) two of the balls drawn are white, (ii) one is of each colour,
(iii) none is red, (iv) at least one is white.
1
Expert's answer
2020-03-10T13:13:56-0400

There are 19 balls of which 6 are white, 4 are red and 9 are black. Number of ways N 3 balls can be drawn is 


"\binom{19}{3}={19! \over 3!(19-3)!}={19(18)(17) \over 1(2)(3)}=969"


(i) Find the probability that two of the balls drawn are white


"P(two\ white)={\binom{6}{2}\binom{19-6}{1} \over 969}={{6! \over 2!(6-2)!}(13) \over 969}={195 \over 969}={65 \over 323}"

(ii) Find the probability that one is of each colour


"P(each\ color)={\binom{6}{1}\binom{4}{1}\binom{9}{1} \over 969}={6(4)(9) \over 969}={216 \over 969}={72 \over 323}"

(iii) Find the probability that none is red

Out of 19 balls 4 are red and 15 are not red.


"P(no \ red)={\binom{19-4}{3} \over 969}={15! \over 3!(15-3)!}\cdot{1 \over 969}=""={15(14)(13) \over 1(2)(3)(969)}={455\over 969}"

(iv) Find the probability at least one is white.

Out of 19 balls 6 are white and 13 are not white.


"P(no \ white)={\binom{19-6}{3} \over 969}={13! \over 3!(13-3)!}\cdot{1 \over 969}=""={13(12)(11) \over 1(2)(3)(969)}={286\over 969}"

"P(at\ least\ one \ white)=1-P(no \ white)=""=1-{286\over 969}={683\over 969}"